ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 258, 2023

Artebani, M., C. Correa Deisler and X. Roulleau Mori dream K3 surfaces of Picard number four: projective models and Cox rings

81

Correa Deisler, C. See M. Artebani Doran, C. F. and J. Kostiuk Geometric variations of local systems

and elliptic surfaces

1

Glutsyuk, A. Density of a thin film billiard reflection pseudogroup in a Hamiltonian symplectomorphism pseudogroup

137

Haulmark, M. and G. C. Hruska On canonical splittings of relatively hyperbolic groups

249

Hruska, G. C. See M. Haulmark Kostiuk, J. See C. F. Doran Kuhn, M. G., S. Saliani and T. Steger Free group representations

from vector-valued multiplicative functions. III

339

Lian, Y. See D. Wu Liang, Z. and Z. Wang Reducibility of 1-d Schrödinger equation

with unbounded oscillation perturbations

287

Mondal, A. n-Kazhdan groups and higher spectral expanders  453

Roulleau, X. See M. Artebani Saliani, S. See M. G. Kuhn Siciliano, G. and K. Silva On the structure of the Nehari set asso-

ciated to a Schrödinger–Poisson system with prescribed mass: old and new results

403

Silva, K. See G. Siciliano

Spiegelhofer, L. Collisions of digit sums in bases 2 and 3  475

Steger, T. See M. G. Kuhn Wan, C. and L. Zhang The multiplicity problems for the unitary

Ginzburg–Rallis models

185

Wang, Z. See Z. Liang Wu, D., Y. Lian and K. Zhang Pointwise boundary differentiability

for fully nonlinear elliptic equations

375

Zhang, K. See D. Wu Zhang, L. See C. Wan